### v3

# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

Abstract
In-context learning (ICL) has emerged as an effective solution for few-shot learning with large language models (LLMs). However, how LLMs leverage demonstrations to specify a task and learn a corresponding computational function through ICL is underexplored. Drawing from the way humans learn from content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword, and template tokens. Our goal is to identify the types of tokens whose representations directly influence LLM’s performance, a property we refer to as being performance-critical. By ablating representations from the attention of the test example, we find that the representations of informative content tokens have less influence on performance compared to template and stopword tokens, which contrasts with the human attention to informative words. We give evidence that the representations of performance-critical tokens aggregate information from the content tokens. Moreover, we demonstrate experimentally that lexical meaning, repetition, and structural cues are the main distinguishing characteristics of these tokens. Our work sheds light on how large language models learn to perform tasks from demonstrations and deepens our understanding of the roles different types of tokens play in large language models.
Identifying and Analyzing Performance-Critical Tokens in
Large Language Models
Yu Bai1††thanks: Work in progress., Heyan Huang1, Cesare Spinoso-Di Piano2,3, Marc-Antoine Rondeau2, Sanxing Chen4, Yang Gao1 ††thanks: Corresponding author., and Jackie Chi Kit Cheung2,3,5 1Beijing Institute of Technology 2Mila – Quebec Artificial Intelligence Institute 3McGill University, 4Duke University 5Canada CIFAR AI Chair yubai@bit.edu.cn
1 Introduction
In-context learning (ICL) has become a popular technique employed with large language models (LLMs) (Brown et al., 2020). However, ICL has been shown to be unstable in that slight changes to the in-context prompts (e.g., reordering of demonstrations) can lead to substantial differences in performance (Lu et al., 2022; Zhang et al., 2022). This circumstance is difficult to control due to a lack of understanding of the model’s working mechanisms, leaving us uncertain about the exact process by which LLMs learn to infer a task specification from demonstrations and produce a computation function to implement that task specification. Previous papers have begun to explore this issue, focusing on specific aspects such as the label space (Min et al., 2022) and the hidden states of the last prompt token (Hendel et al., 2023; Todd et al., 2023), but have been limited in scope.
In this work, we conduct a comprehensive study on how LLMs extract information that is valuable for improving task performance from demonstrations. Drawing from the way humans learn through content-label mappings in demonstrations, we categorize the tokens in an ICL prompt into content, stopword (Wilbur and Sirotkin, 1992), and template tokens, with the latter two typically ignored by humans due to their uninformative nature (Lenartowicz et al., 2014). We use content tokens as a category because they “are fixated (by human) about 85% of the time” (Rayner, 1998). With these categories in mind, we ablate the representations of different token types from the attention of ICL test examples, masking partial information during the model’s task-solving process, as shown in Figure 1. This ablation is intended to identify the types of tokens whose representations LLMs directly depend on to achieve high-level performance, thereby explaining how LLMs learn from demonstrations. We refer to these tokens that are critical for performance as performance-critical tokens.
We show that template and stopword tokens (e.g., “Answer:”) are the most prone to be performance-critical tokens. In contrast, content tokens (e.g., “Union”, “player”, etc.) have a negligible impact on performance when their representations are eliminated from the attention of the test examples. This finding is counterintuitive since the original text of template and stopword tokens inherently do not possess any information found in the demonstrations. To explain this, we study the relationship among different types of tokens through ablation experiments that cut off the information flow between different kinds of tokens. We show that content tokens are indirectly leveraged by LLMs during ICL through aggregating their information into the representations of performance-critical tokens, raising questions about how LLMs use these tokens in general text processing and reasoning.
Beyond identifying performance-critical tokens, we analyze them to better understand how they are leveraged by LLMs. Specifically, we investigate the characteristics which differentiate them from other tokens. We find the following three distinguishing characteristics: the lexical meaning of tokens as it relates to the task being solved, the repetition of tokens throughout the prompt, and the structural cues which the tokens provide to the prompt. Our findings indicate that the lexical meaning, repetition, and structural cues of performance-critical tokens contribute to task performance across all model sizes, suggesting that these characteristics contribute to the property of tokens being performance-critical.
Our work reveals that we can identify and characterize the types of tokens whose representations are the most important in directly maintaining ICL task performance. Our results suggest that previous claims about ICL should be more nuanced, in that representations of tokens beyond label words (Wang et al., 2023) may also directly impact the task performance. We investigate the characteristics of lexical meaning, repetition, and structural cues related to performance-critical tokens which allow us to partially explain their importance and help us better understand how to avoid performance instability while using ICL. Additionally, our findings reinforce that we cannot assume LLMs and humans solve tasks in similar ways, emphasizing the need to develop better explanations of LLM behaviors. Overall, our research deepens the understanding of the roles different types of tokens play in LLMs, pointing to future work that explores how specific token representations can be leveraged for particular purposes (e.g., storing compressed information). Code and data will be released in the camera-ready version.
2 Related Work
2.1 Working mechanisms of ICL
Since the proposal of in-context learning (Brown et al., 2020), its working mechanisms have been extensively studied by the research community (Min et al., 2022; Liu et al., 2021; Olsson et al., 2022; Bhattamishra et al., 2023; Zhou et al., 2024). Min et al. (2022) suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, Yoo et al. (2022) propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. Mao et al. (2024) analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (Han et al., 2023). Zhao et al. (2024) propose to use coordinate systems to understand the working mechanism of in-context learning. Zhou et al. (2023) propose a comprehensive survey on the interpretation and analysis of in-context learning.
Our work investigates the working process of ICL in LLMs at inference time, demonstrating that certain specific tokens are more likely to possess representations that could affect the processing of final test samples, improving the task performance.
2.2 Function vectors of in-context learning
Todd et al. (2023) and Hendel et al. (2023) provide evidence that function vectors store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt, which can then be added to or replace the corresponding vectors in a zero-shot example, yielding results comparable to using all demonstrations as context. Liu et al. (2023a) propose using an in-context vector to represent the target task and apply feature shifting to query examples. They feed each input and its target into an LLM, concatenate the latent states, and apply PCA to derive a vector aligned with the task. Finally, Wang et al. (2023) propose that label words in demonstrations act as information anchors, aggregating information from previous examples to provide it to the test example. This suggests label tokens may satisfy our definition of performance-critical tokens.
All these previous studies either solely focus on a single token (i.e., the last prediction prompt token or label token) of the ICL prompt or treat the entire demonstration as a single unit, neglecting the other tokens within it. Our research focuses on all the tokens in the prompt and reveals that there are additional tokens with specific characteristics whose representations significantly affect the final ICL performance.
3 Preliminaries
3.1 Notation
In in-context learning (ICL), task demonstrations (e.g., input-output pairs) are leveraged to construct a structured prompt that guides the model in predicting the final answer. Formally, the structural prompt consists of the following components: the instruction , the templates , , and the demonstrations , , where denotes the th demonstration while and refer to the input text and output labels, respectively. These prompt components are concatenated to form the ICL prompt, , as shown in Table 1. During inference, the templated version of the test example without its answer, , is appended to the ICL prompt and then sent to the large language model to predict the corresponding answer. We use to denote the concatenation of token sequences.
Component notation | Component example |
[…]
Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |
Business | |
Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |
Sports | |
ICL Prompt | Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: Radio veteran Karmazin joins Sirius. Sirius Satellite Radio Inc. named former Viacom Inc. president Mel… | |
Answer: Business | |
Article: Numbers point to NY. NEW YORK - The New York Yankees can achieve two milestones with one more victory… | |
Answer: Sports |
3.2 Definition of performance-critical token
In defining the performance-critical tokens, we measure the performance variation before and after incorporating the representations of specific tokens into the attention scope of the test example. Specifically, we define performance-critical tokens as the tokens that lead to both a noticeable performance improvement when their representations are included in the attention of test examples and performance degradation when they are excluded.
Formally, let be a large language model and be a classification dataset. We define as the set of representations of each token in the ICL prompt and as the set of representations of the test demonstration which is appended to for prediction (i.e., ). In addition, we let be some set of representations which may attend to from at inference time while performing ICL. For instance, would imply that, when is predicting the label of the test demonstration, the attention from the test example is restricted to the prompt’s instruction token representations.
Then, in order to provide a practical definition for the performance-critical tokens, we let be the accuracy achieved by a LLM when performing ICL on the classification dataset where the only representations which the test example may attend to at inference time are . Given a partition of , we say that a set of tokens is performance-critical if
We note that examining the possibility of each token being performance-critical (i.e., ) in an ICL prompt would be computationally intractable. We instead categorize all the tokens based on the role they play in the prompt and identify which types of tokens are more likely to be performance-critical in Section 5.
4 Experimental settings
Datasets
we consider the most widely used text classification datasets used by previous studies (Zhao et al., 2021). For topic classification, we use the 4-way and 14-way datasets AGNews and DBPedia (Zhang et al., 2015). For textual entailment, we use the 3-way CB (De Marneffe et al., 2019) and 2-way RTE dataset (Dagan et al., 2005). We also use SST2 (Socher et al., 2013) and TREC (Voorhees and Tice, 2000) for sentiment and question classification tasks. Besides these classification tasks, we also present results in machine translation and question answering tasks to show that our findings can also be extended to text generation tasks. Results and analyses of these tasks are attached to Appendix L and Appendix M.
Evaluation
For each dataset, we randomly select 4 training demonstrations from the training set using 15 seeds, limited by the computational cost of LLM inference. For testing, we evaluate each setting on 500 randomly selected test examples. We show this sample size is sufficient by comparing results with 500 test examples and the full dataset using OpenLlama 3B and Llama 7B models, as shown in Appendix J. Instruction prompt is retained in all ablations, as it is essential for model performance (Yin et al., 2023). We use a fixed for all main results, with additional experiments using different provided in Appendix K, demonstrating that changing does not affect our main findings.
LLMs
We utilize the 7B, 13B, and 33B Llama models and a 3B OpenLlama model. Additional results with Llama 2 7B, Llama 2 13B, and Mistral 7B models are demonstrated in Appendix F, and models after supervised fine-tuning are tested in Appendix G. All experiments are conducted on a single A100 80G GPU. For the 13B and 33B models, we apply 8-bit quantization to fit them into a single GPU. The experiments are conducted using Huggingface Transformers (Wolf et al., 2020).
5 Performance-Critical Token Detection
In this section, we aim to identify the performance-critical tokens in the ICL prompt. We first structurally categorize all the tokens in the prompt into three types: template, stopword, and content tokens. Then, we provide supporting evidence from the view of task performance to show that the template and stopword tokens are the most prone to be performance-critical tokens. Finally, we demonstrate that the information of content tokens serve to indirectly contribute to the performance by being propagated into the representations of the performance-critical tokens by LLMs.
5.1 Token types
We categorize ICL tokens based on the structure of the ICL prompt, following our notation in Table 1. Firstly, we find it natural to categorize tokens based on the structure of ICL prompts where the tokens from the demonstration examples and the labels are separated by template tokens from and . Second, can be subdivided into content and stopword tokens, with the latter typically providing less useful information and often being ignored (Rayner, 1998) when humans use analogy to learn specific tasks. Guided by these intuitions, we categorize all the tokens in the ICL prompt into template tokens, stopword tokens, and content tokens as follows:
Template tokens (temp): In defining template tokens, we include all the tokens which serve as templates for the ICL prompt. This includes the tokens in and , as shown in Table 1.
Stopword tokens (stop): In defining stopword tokens, we include punctuation and conjunction words, such as [,], [.], etc., in the prompt. We use the stopword tokens appearing in the instructions111The stopword token list used in the main experiments and the ablation results with the complete NLTK (Loper and Bird, 2002) stopwords list are shown in Appendix H..
Content tokens (cont): In defining content tokens, we include all the tokens from except for the ones that are already stopword tokens. We use the term “content tokens” as they convey the meaningful information found in the demonstrations.
Researchers might typically expect content tokens to be critical, as they account for most of the information in the demonstrations. However, in the following experiments, we find that the representations of template and stopword tokens have the greatest direct impact on performance.
5.2 Ablation on token types
To determine which token types are more likely to be performance-critical tokens whose representations directly affect the final performance significantly, we design two experiments which ablate representations or tokens based on token types. The first involves keeping and masking representations of different token types from the attention of the test example. The second involves dropping the various kinds of tokens from the ICL prompt. The main purpose of the first experiment is to identify the performance-critical tokens defined in Section 3.2, while the second experiment aims to cut off the information propagation of different types of tokens to further explore the functioning of performance-critical tokens. Illustrations of these two methods which we refer to as representation-level and token-level ablations are shown in Figure 1 and Figure 2. Examples for the representation-level ablation is provided in Appendix E.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 | 19.5 |
cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 14.8 | |
[…]
temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | 37.4 | |
Llama 7B | Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 | 25.6 |
cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | 15.6 | |
stop | 57.3 | 83.7 | 49.8 | 43.0 | 55.9 | 50.7 | 31.1 | |
temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | 45.3 | |
Llama 13B | Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 | 19.0 |
cont | 27.7 | 52.4 | 33.5 | 10.9 | 61.7 | 41.7 | 19.0 | |
stop | 72.2 | 73.5 | 46.8 | 50.7 | 58.6 | 30.6 | 36.4 | |
temp | 80.0 | 92.3 | 58.6 | 76.9 | 68.5 | 47.7 | 51.7 | |
Llama 33B | Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 | 54.6 |
cont | 24.4 | 61.7 | 62.1 | 10.5 | 65.2 | 63.6 | 6.7 | |
[…]
temp | 80.5 | 95.2 | 65.2 | 75.2 | 79.0 | 80.0 | 24.6 |
5.2.1 Representation-level ablation
Our first ablation stems from the intuition that if LLMs rely on the representations of specific token types to achieve high-level performance, they should perform adequately with only these representations. Performance should drop significantly if these tokens are removed from the test example’s attention. Hence, we first pass the entire ICL prompt to the LLM and then restrict the test example’s attention to representations of a particular token type (or types)222Since tokens have been shown to significantly impact performance (Wang et al., 2023), we always preserve the attention on the representations of the tokens. during its solving of the task. We compute task performances with every possible ablation combination, removing the representations of one (e.g., Standard ICL temp) or two token types (e.g., Zero-shot cont333Removing two types of tokens from Standard ICL is equivalent to adding the other type to Zero-shot.) from the attention of the test example. All the task performances and the averaged relative performance changes are reported, shown in Table 2 and 3. An illustration of this set of experiments is shown in Figure 1.
[…]
Rare exceptional cases appear when performance is relatively poor with Standard ICL (e.g., OpenLlama 3B in TREC). In some cases, masking the representations of the content tokens brings even better performance than the Standard ICL method, which is possibly due to the elimination of noisy information in the demonstration content. Another interesting observation is that the performance results of Standard ICLstop and Standard ICLcont where the attention to the content and stopword tokens is ablated respectively are similar, with an average difference of only 5.4%. This indicates that the representation of stopword tokens may contain overlapping information with their preceding content tokens. We believe that this could enable LLMs to model long sequences without significant architectural changes (e.g., using stopword representations as synthesis checkpoints) and leave the verification of this hypothesis to future work.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
cont | 58.2 | 86.9 | 27.6 | 61.9 | 56.5 | 51.7 | 0.9 | |
stop | 51.8 | 78.9 | 28.8 | 30.3 | 53.6 | 45.2 | 9.9 | |
temp | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | 23.8 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
cont | 77.9 | 91.5 | 58.5 | 66.5 | 67.8 | 74.4 | 2.0 | |
stop | 78.5 | 88.7 | 39.3 | 66.7 | 60.6 | 60.4 | 9.1 | |
temp | 20.8 | 58.2 | 32.4 | 11.6 | 54.4 | 46.0 | 37.6 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
cont | 81.4 | 93.1 | 58.9 | 75.7 | 69.6 | 45.1 | 0.2 | |
stop | 79.8 | 85.8 | 64.4 | 73.6 | 64.5 | 40.2 | 2.4 | |
temp | 27.8 | 52.4 | 33.5 | 10.9 | 63.1 | 45.6 | 31.5 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
cont | 82.3 | 95.4 | 64.9 | 76.1 | 80.4 | 82.0 | 1.5 | |
stop | 84.8 | 94.9 | 62.1 | 77.3 | 70.5 | 74.4 | 4.3 | |
temp | 24.4 | 61.7 | 60.6 | 10.5 | 67.7 | 68.5 | 32.7 |
5.2.2 Token-level ablation
In this section, we modify the ICL prompt by removing certain types of tokens from the ICL prompt555For template tokens, this includes both the tokens in the demonstrations and the test example to maintain their consistency. We included the analyses of only ablating the tokens in the demonstrations in Appendix P. to further investigate the relationship between different kinds of tokens, cutting off the information flow between the representations of different tokens, shown in Figure 2. When we ablate the template tokens, we preserve the answer and next-line tokens in the templates to maintain a basic separator between the demonstration inputs and outputs. Results averaged on all the datasets are presented in Figure 3. Detailed results on each dataset could be seen in Appendix O.
Our first finding from this ablation is that information is propagated from content token representations to performance-critical token representations, as shown by the contrast between representation-level and token-level ablation in Table 4. The representations of template and stopword tokens alone (i.e., Standard ICL cont in Figure 3) are less effective at affecting task performance, leading to worse performance than those where content token information is included in their attention (i.e., Standard ICL cont in Table 3).
This finding provide us with additional insights about how LLMs leverage different kinds of tokens during ICL. Firstly, this circumstance means that even though the representations of the content tokens are not directly used when LLMs predict the answer, the encoding of these tokens contribute to the final performance indirectly through being aggregated into the representations of the performance-critical tokens. Secondly, it also suggests that LLMs prefer to utilize the the performance-critical tokens to aggregate the indirect information from the demonstration rather than others (i.e., content tokens). It is their incorporation of this information that makes them better at encoding tasks, partially explaining the working mechanism of ICL.
Ablation( cont) | OpenLlama 3B | Llama 7B | Llama 13B | Llama 33B |
Rep. level | 57.1 | 72.8 | 70.6 | 80.1 |
Token level | 44.3 | 54.9 | 65.1 | 70.4 |
Difference | 12.8 | 17.9 | 5.5 | 9.7 |
In addition, removing template tokens causes the LLMs to completely lose their ability to solve tasks via ICL with an overall task accuracy performance of 0% for all sizes and all tasks. We hypothesize that this is because the model no longer has an explicit cue to generate the target label, which is further discussed in Section 6. In this case, if we add back the last prompt token after the next-line token, the results return to their original level due to the introduction of a template token. This finding confirms previous claims that preserving the format of ICL prompts plays a significant role in retaining the task performance (Min et al., 2022).
Roles of different types of tokens
To summarize, template and stopword tokens are the most performance-critical. Template token representations significantly improve the task performance, while the representations of stopword tokens play a more supportive role in the spectrum of performance-critical tokens by summarizing the information of content tokens. In contrast, content token representations do not directly contribute to the performance but instead indirectly provide information that is aggregated into the other two types of tokens. We discuss possible applications of these findings in Appendix Q. This finding further raises an additional question: What are the characteristics for a token to be perceived by a LLM as performance-critical?
6 Performance-Critical Token Analyses
To answer the above question, we provide analyses of the tokens whose representations we believe mainly store information that directly affects the performance of a task drastically. We focus on the template tokens since, as evidenced by the findings in Section 5.2.1 and 5.2.2, their representations are the most important to maintaining task performance. Our analysis focuses on the distinguishing characteristics of performance-critical tokens, while we also examine the effects of each part of template tokens in Appendix C.
Settings | Notations | Examples |
[…]
By better understanding what characteristics of performance-critical tokens lead them to affect task performance, we provide insights on how to best leverage LLMs for ICL (e.g., What principles should practitioners be using when designing prompt templates?). We hypothesize that the following characteristics are critical for a token to be leveraged as performance-critical tokens: lexical meaning referring to the task-related lexical meaning of a performance-critical token, repetition referring to the multiple appearances of the performance-critical tokens in the prompt, and structural cues referring to how performance-critical tokens format the ICL prompt, shown in Table 1, into structured text.
We design several experiments to test if these characteristics affect the impact of performance-critical tokens, by disrupting each characteristic in the ICL prompts. A characteristic is related if there is a performance drop after the disruption. The disruption is achieved by replacing the template tokens with different kinds of random string templates, shown in Table 5. We use 5 different random string templates which are attached to Appendix T and average all the results for each setting.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
Swap | 64.4 | 86.8 | 21.7 | 58.7 | 60.6 | 54.6 | 57.8 | |
Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
Swap | 70.2 | 11.4 | 44.3 | 58.2 | 64.5 | 50.1 | 49.8 | |
Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
Swap | 81.5 | 67.4 | 36.4 | 75.9 | 69.1 | 52.1 | 63.7 | |
Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
Swap | 84.5 | 94.9 | 60.8 | 75.5 | 68.0 | 55.5 | 73.2 | |
Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 |
Lexical meaning
A performance-critical token might be more impactful on the performance based on its lexical meaning. One possible hypothesis is that if the token carries specific task-related meanings like “Answer”, it is more likely to serve as a performance-critical token.
To verify if lexical meanings could affect the formation of performance-critical tokens, we 1) Replace the tokens from and with the same random strings across the different demonstrations (Randomfixed), thus completely disrupting the lexical characteristic of these tokens; 2) Swap and (Swap), thus partially disrupting the lexical characteristic of these tokens. Table 6 shows that disrupting the lexical meaning of tokens slightly impacts task performance in smaller models (OpenLlama 3B), while larger models experience more significant drops. Llama 7B, in particular, is highly sensitive to lexical meaning and performs worse when semantics are disturbed. Thus, the lexical meaning of tokens likely influences their performance-critical nature.
Repetition
The impact of performance-critical tokens could also be influenced by their repetition throughout the prompt. Intuitively, via the attention mechanism, repetitive patterns are more likely to propagate information through the processing of text. Yan et al. (2024) propose self-reinforcement in in-context learning, also suggesting that repetition could be a significant factor in ICL.
We explore the repetition characteristic by comparing the results of the previously discussed Randomfixed experiment with an experiment replacing and with different random strings (Randomnonfixed), thus breaking the repetition of template tokens present in ICL demonstrations. We further conduct supplemental experiments using template tokens with specific lexical meanings for comparison, as detailed in Appendix U.
Table 7 shows that without consistent repetition of performance-critical tokens, performance decreases for most models. This suggests that necessary information may not have been properly accumulated in the template token representations. These experiments demonstrate that repetition significantly influences the impact of performance-critical tokens. The results align with previous findings, reinforcing our claim that repetition is a key characteristic of performance-critical tokens.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 |
Randomnonfixed | 30.2 | 71.4 | 17.1 | 18.6 | 47.9 | 47.7 | 38.8 | |
Llama 7B | Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 |
Randomnonfixed | 15.5 | 11.6 | 10.4 | 1.8 | 4.6 | 25.6 | 11.6 | |
Llama 13B | Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 |
Randomnonfixed | 32.1 | 34.5 | 19.2 | 6.0 | 21.0 | 32.8 | 24.3 | |
Llama 33B | Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 |
Randomnonfixed | 78.5 | 87.5 | 46.3 | 63.1 | 63.6 | 46.1 | 64.2 |
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 70.7 | 51.7 | 40.4 | 53.5 | 50.2 | 48.6 | 53.3 |
Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 | |
Llama 7B | Standard ICL | 72.3 | 77.4 | 54.1 | 64.7 | 53.0 | 64.4 | 64.3 |
Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 | |
Llama 13B | Standard ICL | 82.0 | 72.0 | 60.1 | 75.9 | 60.4 | 18.8 | 70.1 |
Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 | |
Llama 33B | Standard ICL | 85.3 | 88.3 | 71.2 | 75.5 | 64.1 | 45.5 | 76.9 |
Randomfixed | 69.7 | 53.0 | 37.8 | 72.8 | 53.0 | 37.6 | 54.0 |
Structural cues
Beyond lexical meaning and repetition, the influence of performance-critical tokens may also depend on how ICL prompts are formatted. ICL prompts often include structural cues to assist the model to differentiate between elements with distinct roles, such as task inputs and target labels, within a demonstration. For example, template tokens (i.e., and ) delimit demonstration examples and labels, while stopword tokens (e.g., ,”, .”, “:”, etc.) structure content words into sentence components. Examples of how performance-critical tokens delimit ICL prompts are shown in Appendix W. These structural cues are similar to those in LLM pretraining data (e.g., column names in SQL tables), suggesting that pretraining on such data enables the model to recognize the structuring role of performance-critical tokens, allowing their representations to store higher-level information.
To assess the structuring characteristic of performance-critical tokens, we perturb the structure of one-shot ICL prompts in two stages, where the one-shot setting could eliminate repetition as a confounding factor. First, we disrupt the lexical meaning of template tokens, since token meaning helps LLMs distinguishing the different parts of a prompt. Then, we remove all template tokens to eliminate any source of structure cues. Table 8 shows that disrupting structural cues decreases performance, highlighting their importance. Consistent with Section 5.2.2, removing all template tokens results in 0% performance due to the complete elimination of structural cues. Supplemental experiments in Appendix V further support this from a representation-level ablation perspective.
7 Conclusion
In this paper, we have provided a fine-grained characterization of performance-critical tokens, whose representations LLMs directly depend on to achieve high-level performance in ICL. Through a series of experiments, we have examined the roles of template tokens and stopword tokens within ICL as potential performance-critical tokens. Our findings add nuance to previous claims made about ICL, for example, that tokens other than label words could also provide valuable information directly affecting the performance. Overall, our results demonstrate that model performance depends directly on the presence of these tokens and that their lexical meaning, their repetition throughout the ICL prompt, and their structural formatting of ICL demonstrations are likely to play a role in how effectively they allow an LLM to recover the critical information needed to perform a task.
Limitations
In this paper, the token categorization is performed manually, leaving room for further refinement, leaving the exploration of other specific content tokens as performance-critical tokens in certain contexts to future work. While the results provide robust support to our categorization, the identification process itself lacks precision. For instance, stopwords may only represent a subset of all in-context performance-critical tokens. The manual nature of our categorization limits our ability to comprehensively track these tokens. Moreover, our experiments are limited to classification, machine translation, and question answewring datasets, suggesting that our conclusions should be further validated for other tasks. Additionally, our focus on performance-critical tokens, whose representations could impact task performance, may overlook other tokens responsible for other possible functions. Another limitation of our study is that we focus exclusively on in-context learning scenarios, meaning that our findings may not be directly applicable to zero-shot learning scenarios.
Ethics Statement
This work focuses on analyzing the working mechanisms of large language models and, as such, does not present any increased risks of harm beyond the existing norms of natural language processing or computational linguistics research. The associated risks include using a model trained on vast amounts of text, which may inadvertently contain biases. Another concern is the potential misuse of the model for generating misleading or harmful content. However, such a scenario is unlikely in our work, as we concentrate on classification tasks with fixed outputs.
Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we have made several efforts that are documented throughout this paper. Our experiments utilize the open-source models described in Section 4. The prompts and templates used in our experiments are detailed in Section 3.1, Section 6 of the main text and in Appendix D, Appendix E, Appendix K, Appendix T, Appendix U. The stopword token list used in our experiments is shown in Appendix H. The complete code for our implementation, including all inference processes, is provided in the supplementary materials. We employed random seeds ranging from 1 to 15 to ensure consistent results across experiments, as specified in Section 4 and the supplementary code. All datasets used in our experiments are described comprehensively in Section 4, and the supplementary code includes all data processing steps and any preprocessing applied. We encourage other researchers to consult these references for replicating our findings.
Datasets | Notations | Examples |
[…]
Datasets | Stopwords |
[…]
NLTK Stopwords List |
[…]
Models | Settings | RTE | Settings | RTE | ||
with “:” | w/o “:” | with “:” | w/o “:” | |||
Llama 7B | temp with | 66.3 | 59.5 | temp w/o | 40.7 | 42.5 |
58.9 | 56.5 | 43.7 | 49.9 | |||
56.7 | 56.7 | 56.0 | 55.6 | |||
Llama 13B | temp with | 68.5 | 59.8 | temp w/o | 57.5 | 53.7 |
65.5 | 59.0 | 53.6 | 52.8 | |||
61.2 | 58.4 | 54.8 | 53.7 | |||
Llama 33B | temp with | 79.0 | 77.1 | temp w/o | 71.8 | 65.8 |
77.4 | 75.4 | 70.6 | 67.8 | |||
72.8 | 70.0 | 67.0 | 61.3 |
Appendix B More Related Work
B.1 Working mechanisms of ICL
Since the proposal of in-context learning (Brown et al., 2020), its working mechanisms have been extensively studied by the research community (Min et al., 2022; Liu et al., 2021; Olsson et al., 2022; Bhattamishra et al., 2023). Min et al. (2022) suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, Yoo et al. (2022) propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. Xie et al. (2021) explain ICL as implicit Bayesian inference, while Akyürek et al. (2022) explore ICL learning process using linear models. Theoretical explanations (Guo et al., 2023; Bai et al., 2023; Li et al., 2023b) and gradient descent explanations have also been proposed. Mao et al. (2024) analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (Han et al., 2023). Zhao et al. (2024) propose to use coordinate systems to understand the working mechanism of in-context learning. Zhou et al. (2023) propose a comprehensive survey on the interpretation and analysis of in-context learning.
Additional analyses exploring different aspects of ICL have also been studied. For instance, order sensitivity where task performance fluctuates based on the order of the same ICL demonstrations has been identified as a limitation of ICL (Lu et al., 2022). Yan et al. (2024) propose that repetitive patterns in the prompt could affect the ICL performance in both positive and negative ways. Pan et al. (2023) analyze the ICL process by disentangling it into task recognition and task learning. Madaan and Yazdanbakhsh (2022) propose to define text and patterns while using counterfactual prompting for attributing token importance in chain-of-thought techniques.
Our work investigates the working process of ICL in LLMs at inference time, demonstrating that certain specific tokens are more likely to possess representations that could affect the processing of the final test sample, improving the task performance.
B.2 Function vectors of in-context learning
Todd et al. (2023) and Hendel et al. (2023) provide evidence of function vectors that store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt. These vectors can then be added to, or used to replace, the corresponding vectors in a zero-shot example, achieving results comparable to those obtained when the model uses all demonstrations as context. In addition, Liu et al. (2023a) also propose using an in-context vector to represent the target task and applying feature shifting to query examples. They first feed each input and its corresponding target separately into an LLM, then concatenate all the latent states. A PCA method is applied to derive a vector that is more closely aligned with the task. Finally, Wang et al. (2023) propose that label words in the demonstration examples function as information anchors by aggregating the information from previous demonstrations and providing it to the test example. This finding suggests that we may view label tokens as satisfying our definition of performance-critical tokens.
All these previous studies either solely focus on a single token (i.e., the last prediction prompt token or label token) of the ICL prompt or treat the entire demonstration as a single unit, neglecting the other tokens within it. Our research focuses on all the tokens in the prompt and reveals that there are additional tokens with specific characteristics whose representations significantly affect the final ICL performance.
Appendix C Effects of different performance-critical tokens
In this section, we aim at examining the relationship among the representations of different performance-critical tokens. To achieve this, we test the effectiveness (i.e., how much they could affect the downstream task performance) of each part of performance-critical tokens to see if they could work without each other.
To achieve this, we ablate the representations of each performance-critical token, similar to Section 5.2.1. In Section 5.2.1, we assume that the label token is needed for ICL to achieve performance results on par with Standard ICL, as suggested by previous work (Wang et al., 2023). However, it is still not known how the other performance-critical tokens affect the performance without . Hence, we divide our experiments by including or excluding the label tokens to further specifically investigate their effectiveness. We present the results on RTE datasets in Table 12 while full results are shown in Appendix R.
Overall, the above experiments show that the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration or instability issues. From the results with , it is observed that all the template tokens (i.e., , , and “:”) contribute to the final performance. Removing one of them would cause a performance degradation. From the results without , the performance becomes less predictable, where adding back a template token (e.g., “:”) does not always bring performance improvements. Moreover, in some datasets, models without can still achieve relatively high performance. These results show that representations of other template tokens may also be seen as information anchors whose representations aggregate and serve information to the final prediction of LLMs, broadening the conclusions of Wang et al. (2023) who claim that only answer tokens serve as information anchors.
Appendix D In-context Learning Templates
[…]
Dataset | AGNews | DBPedia | SST2 | TREC | RTE | CB |
Test set number of the dataset | 7,601 | 70,000 | 1,821 | 500 | 277 | 250 |
Test set number we scaled up | 5,000 | 5,000 | 1,821 | - | - | - |
OpenLlama 3B | |||||||
Setting | AGNews | SST2 | DBPedia | Setting | AGNews | SST2 | DBPedia |
Zero-shotcont | 25.7 | 52.4 | 7.8 | Standard ICLcont | 57.6 | 86.4 | 62.3 |
Zero-shotstop | 37.2 | 82.6 | 53.1 | Standard ICLstop | 62.0 | 91.0 | 63.2 |
Zero-shottemp | 56.2 | 86.3 | 62.5 | Standard ICLtemp | 41.5 | 87.2 | 57.3 |
Llama 7B | |||||||
Setting | AGNews | SST2 | DBPedia | Setting | AGNews | SST2 | DBPedia |
Zero-shotcont | 32.0 | 58.1 | 13.4 | Standard ICLcont | 77.0 | 91.9 | 67.4 |
Zero-shotstop | 58.0 | 84.6 | 43.9 | Standard ICLstop | 79.5 | 94.2 | 69.0 |
Zero-shottemp | 71.4 | 90.7 | 67.4 | Standard ICLtemp | 64.3 | 84.9 | 58.6 |
Appendix E Representation-level Ablation Examples
We provide a one-shot demonstration of all ablation cases for the representation-level ablation experiments, as shown in Table 15. In these demonstrations, representations of specific token types are masked in the attention mechanism, where <m> denotes that all representations of the token are removed from the attention scope from the test example.
Zero-shottemp |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> |
Answer: Technology |
Zero-shotstop |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> <m> <m> <m> the <m> . <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> the <m> <m> <m> <m> <m> <m> <m> <m> , <m> <m> <m> <m> <m> <m> of <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> into <m> <m> the <m> <m> . |
<m> <m> Technology |
Zero-shotcont |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> First class to <m> moon <m> London - British airline magnate Richard Branson announced a plan on Monday for <m> world’s first commercial space flights <m> saying “thousands” <m> fee-paying astronauts could be sent <m> orbit in <m> near future <m> <m> <m> <m> Technology <m> <m> |
Standard ICLtemp |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
<m> <m> <m> First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. <m> <m> <m> Technology <m> <m> |
Standard ICLstop |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: First class to <m> moon <m> London - British airline magnate Richard Branson announced a plan on Monday for <m> world’s first commercial space flights <m> saying “thousands” <m> fee-paying astronauts could be sent <m> orbit in <m> near future <m> <m> Answer: Technology <m> <m> |
Standard ICLcont |
<s> Classify the news articles into the categories of World, Sports, Business, and Technology. |
Article: <m> <m> <m> the <m> . <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> the <m> <m> <m> <m> <m> <m> <m> <m> , <m> <m> <m> <m> <m> <m> of <m> <m> <m> <m> <m> <m> <m> <m> <m> <m> into <m> <m> the <m> <m> . |
Answer: Technology |
Appendix F Experiment Results with More Large Language Models
Experimental results using the Llama 2 and Mistral models are shown in Table 16. The trends observed in these experiments are consistent with those involving the Llama models. These results further reinforce the findings of this paper, indicating that template tokens and stopword tokens are the most prone to serving as performance-critical tokens.
Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
Llama 2 7B | Zero-shot | 36.0 | 50.2 | 50.4 | 57.2 | 6.4 | 51.6 | 0.0 |
cont | 1.9 | 0.9 | 61.0 | 50.6 | 12.9 | 48.7 | 53.2 | |
stop | 23.4 | 49.0 | 78.1 | 54.4 | 61.6 | 65.3 | 47.9 | |
temp | 31.3 | 81.1 | 82.6 | 55.2 | 65.5 | 63.9 | 55.4 | |
Llama 2 13B | Zero-shot | 53.3 | 56.2 | 90.8 | 49.0 | 7.6 | 70.0 | 46.4 |
cont | 14.1 | 0.5 | 56.0 | 61.4 | 0.0 | 62.6 | 54.6 | |
stop | 9.6 | 47.2 | 76.8 | 65.2 | 65.3 | 66.5 | 56.8 | |
temp | 18.1 | 78.2 | 93.7 | 62.4 | 70.4 | 71.9 | 52.1 | |
Mistral 7B | Zero-shot | 59.5 | 77.8 | 84.4 | 73.0 | 57.8 | 1.8 | 62.1 |
cont | 10.0 | 43.3 | 52.0 | 66.6 | 10.1 | 64.3 | 60.7 | |
stop | 18.4 | 78.9 | 92.5 | 71.6 | 81.4 | 69.9 | 72.9 | |
temp | 19.7 | 81.7 | 95.9 | 63.9 | 83.3 | 77.9 | 72.5 | |
Llama 2 7B | Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 |
cont | 3.8 | 82.4 | 85.5 | 54.3 | 64.2 | 59.6 | 55.7 | |
stop | 2.5 | 84.8 | 88.0 | 51.7 | 65.7 | 65.8 | 53.2 | |
temp | 32.8 | 0.9 | 61.0 | 50.6 | 12.9 | 48.5 | 53.6 | |
Llama 2 13B | Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 |
cont | 1.3 | 79.0 | 94.1 | 62.7 | 72.4 | 72.1 | 53.6 | |
stop | 2.9 | 80.1 | 89.4 | 61.5 | 74.1 | 69.6 | 49.3 | |
temp | 33.5 | 0.5 | 56.0 | 61.4 | 0.0 | 68.2 | 54.3 | |
Mistral 7B | Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 |
cont | 0.6 | 81.8 | 96.2 | 64.4 | 83.4 | 78.9 | 72.9 | |
stop | 0.8 | 81.3 | 97.0 | 66.5 | 80.5 | 75.5 | 75.7 | |
temp | 22.8 | 78.6 | 52.0 | 66.6 | 10.1 | 67.1 | 69.6 |
Appendix G Experiment Results with Instruction-tuned Large Language Models
We present the representation-level ablation results of large language models after instruction tuning to confirm that our findings remain consistent. Specifically, we use Llama 2 7B Chat and Llama 2 13B Chat in our experiments. As shown in Table 17, the results align with the findings discussed in Section 5.2.1, with only one exception: the TREC dataset. In this dataset, the input data is structured in a question-answering format (e.g., Who/What/When did …). We hypothesize that, during the supervised fine-tuning process, tokens associated with these question formats may also serve as performance-critical tokens, although they are currently categorized as content tokens in our experiments. Overall, these supplemental results prove further evidence to our findings in the main paper.
Models | Setting | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
Llama 2 7B Chat | Zero-shot | - | - | - | - | - | - | - |
cont | 27.9 | 0.7 | 52.6 | 52.2 | 7.9 | 24.8 | 28.9 | |
stop | 72.9 | 77.1 | 90.6 | 60.2 | 75.4 | 66.7 | 67.5 | |
temp | 75.6 | 80.1 | 92.6 | 62.6 | 76.6 | 69.5 | 72.1 | |
Llama 2 13B Chat | Zero-shot | - | - | - | - | - | - | - |
cont | 38.4 | 0.0 | 55.7 | 67.3 | 0.5 | 63.5 | 43.2 | |
stop | 67.1 | 78.5 | 87.6 | 66.9 | 71.4 | 67.2 | 31.1 | |
temp | 72.3 | 82.0 | 93.7 | 65.6 | 72.3 | 72.3 | 47.9 | |
Llama 2 7B Chat | Standard ICL | - | - | - | - | - | - | - |
cont | 76.1 | 80.7 | 93.1 | 62.9 | 76.8 | 70.7 | 72.5 | |
stop | 76.4 | 81.7 | 94.4 | 61.9 | 74.9 | 71.0 | 74.3 | |
temp | 31.4 | 0.7 | 52.6 | 52.2 | 7.9 | 24.1 | 51.1 | |
Llama 2 13B Chat | Standard ICL | - | - | - | - | - | - | - |
cont | 74.7 | 83.4 | 93.6 | 66.4 | 74.3 | 74.6 | 56.1 | |
stop | 69.8 | 78.9 | 94.1 | 57.7 | 72.8 | 70.1 | 45.4 | |
temp | 38.7 | 0.0 | 55.7 | 67.3 | 0.5 | 65.1 | 43.6 |
Appendix H Stopword Tokens
For the results shown in the main paper, we used the stopword token (Sarica and Luo, 2021) list shown in Table 10. This list only includes the stopword tokens from the task instruction, aiming to minimize their presence. We made this choice under the assumption that task-affecting information should be stored densely in a few tokens. Hence, the number of tokens whose representations affect the final task performance significantly should be small.
Nevertheless, one might be curious about the results if we used a more complete stopword list. In this case, we utilize a more comprehensive stopword token list of NLTK666https://gist.github.com/sebleier/554280 shown in Table 11 and conduct the representation-level ablation once more. The results are presented in Table 18. It can be observed that all the conclusions from Section 5.2.1 are still well established. A few results are different from Table 2 and Table 3 because we masked the representations of the “</s>” token in this set of experiments. We claim that this masking does not impact the main findings of these experiments.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB |
OpenLlama 3B | Zero-shot | 22.0 | 20.0 | 23.6 | 5.4 | 44.4 | 1.8 |
+ cont | 26.2 | 52.1 | 30.1 | 7.4 | 51.9 | 37.9 | |
+ stop | 38.0 | 85.1 | 31.6 | 54.6 | 58.8 | 55.7 | |
+ temp | 56.5 | 86.7 | 27.1 | 62.2 | 56.4 | 52.3 | |
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | |
- temp | 42.1 | 87.2 | 25.9 | 56.3 | 58.3 | 57.4 | |
- cont | 57.1 | 88.4 | 27.1 | 62.6 | 56.8 | 52.4 | |
- stop | 61.6 | 90.7 | 24.8 | 62.2 | 56.7 | 51.9 | |
Llama 7B | Zero-shot | 25.0 | 29.2 | 41.4 | 0.0 | 54.2 | 3.6 |
+ cont | 32.4 | 57.9 | 42.5 | 12.5 | 55.5 | 46.1 | |
+ stop | 59.9 | 85.9 | 51.7 | 28.9 | 56.0 | 52.7 | |
+ temp | 70.8 | 90.2 | 58.4 | 66.2 | 66.3 | 73.5 | |
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | |
- temp | 64.7 | 84.1 | 54.0 | 56.7 | 56.1 | 48.2 | |
- cont | 75.4 | 93.8 | 59.8 | 67.5 | 66.8 | 74.8 | |
- stop | 81.4 | 94.2 | 60.5 | 67.9 | 67.6 | 72.1 | |
Llama 13B | Zero-shot | 59.0 | 18.0 | 37.0 | 0.0 | 0.0 | 0.0 |
+ cont | 30.6 | 52.4 | 43.8 | 13.0 | 60.2 | 45.5 | |
+ stop | 72.7 | 78.7 | 49.2 | 27.4 | 58.5 | 27.1 | |
+ temp | 78.5 | 92.3 | 59.0 | 74.2 | 67.4 | 52.3 | |
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | |
- temp | 71.7 | 80.1 | 56.2 | 8.7 | 56.5 | 29.3 | |
- cont | 79.3 | 93.4 | 60.1 | 74.1 | 68.4 | 47.6 | |
- stop | 79.2 | 94.1 | 59.3 | 73.8 | 68.9 | 44.6 | |
Llama 33B | Zero-shot | 70.2 | 88.6 | 60.6 | 30.2 | 58.1 | 19.6 |
+ cont | 27.8 | 61.7 | 61.9 | 10.8 | 64.2 | 68.1 | |
+ stop | 74.7 | 93.6 | 66.9 | 70.8 | 69.1 | 63.8 | |
+ temp | 80.6 | 95.2 | 63.1 | 71.9 | 78.7 | 84.0 | |
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | |
- temp | 79.5 | 93.8 | 58.5 | 62.8 | 68.0 | 68.0 | |
- cont | 82.7 | 95.9 | 62.9 | 74.1 | 79.6 | 83.1 | |
- stop | 84.4 | 96.1 | 61.8 | 72.8 | 79.4 | 82.1 |
Appendix I Analysis of the Next-line Tokens
In this section, we analyze the next-line token, which is ablated whenever any type of the stopword tokens or template tokens are ablated in the representation-level ablation experiments. We analyze this token by not ablating it when the these types of tokens are ablated. Results presented in Table 19 demonstrate that the next-line token is an important performance-critical token, due to the fact that they improved the performance by a large margin compared to the results in Table 3.
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
[…]
temp | 41.9 | 87.2 | 26.0 | 56.3 | 58.5 | 57.4 | 5.4 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
stop | 80.4 | 94.6 | 61.1 | 68.0 | 67.2 | 72.0 | 0.9 | |
temp | 64.5 | 84.1 | 54.0 | 58.0 | 56.8 | 54.3 | 12.8 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
stop | 81.2 | 94.1 | 59.3 | 76.9 | 69.2 | 40.6 | 0.2 | |
temp | 74.1 | 80.0 | 46.5 | 30.6 | 58.3 | 25.4 | 17.9 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
stop | 84.3 | 95.6 | 65.7 | 77.6 | 78.6 | 81.8 | 1.0 | |
[…]
Llama 2 7B | Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 |
stop | 85.5 | 92.7 | 56.4 | 66.6 | 63.6 | 57.1 | 0.4 | |
temp | 69.8 | 82.8 | 56.3 | 58.8 | 67.5 | 42.9 | 7.7 | |
Llama 2 13B | Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 |
stop | 81.2 | 94.5 | 61.2 | 73.7 | 72.0 | 53.2 | 1.0 | |
temp | 71.1 | 95.6 | 61.0 | 72.4 | 72.9 | 54.3 | 2.4 | |
Mistral 7B | Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 |
stop | 81.2 | 97.3 | 65.5 | 82.0 | 77.6 | 73.9 | 0.6 | |
temp | 78.6 | 89.7 | 67.6 | 79.4 | 70.8 | 72.5 | 3.8 |
Appendix J Comparison Experiments with More Text Examples
In the ideal scenario, our experiments would have been conducted on the full test set. However, in practice, this is infeasible for any of the models studied in our paper due to computational resource constraints. For instance, it took 42 hours for the OpenLlama 3B model to run one round of the representation ablation experiment on the whole test set of DBPedia (i.e., one cell in Table 2 and 3 for the DBPedia column). To verify our number of test examples decision, we provide additional results where we scale up the number of test examples and observe no difference with our original experimental setup. Thus, we believe that limiting our test set sample size to 500 is a reasonable setup. We provide the test set statistics and the experiment results in Table 13 and Table 14.
For TREC, RTE, and CB, using 500 test examples won’t affect the final results at all since their test set size is smaller than 500. We provide the results of experiments using all test examples in SST2, and 5000 test examples in AGNews and DBPedia here to prove our point that limiting our test set sample size to 500 is a reasonable compromise. Shown in Table 14, compared to the results we show in Table 2 and Table 3, the numbers are changed less than 1% for all the results.
Appendix K Results using different instruction prompts
We conducted experiments on AGNews and DBPedia with 3 other different instructions to show that the and show the results in Table 21 and Table 22. Based on these additional results, our conclusions remain the same, which shows that our findings are not sensitive to variations of the instruction prompt. The different instruction prompts we used are shown in Table 20.
Datasets | Stopwords |
AGNews Ins. 1 | Classify the text into World, Sports, Business, and Technology. |
AGNews Ins. 2 | Classify the articles based on whether they are in the categories of World, Sports, Business, and Technology. |
AGNews Ins. 3 | Classify the news to World, Sports, Business, and Technology. |
DBPedia Ins. 1 | Classify the text into Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, and Book. |
DBPedia Ins. 2 | Classify the documents into the categories of Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, and Book. |
DBPedia Ins. 3 | Classify the articles based on whether they are in the categories of Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, and Book. |
OpenLlama 3B | |||||||
Setting | Ins. 1 | Ins. 2 | Ins. 3 | Setting | Ins. 1 | Ins. 2 | Ins. 3 |
Zero-shotcont | 26.5 | 26.9 | 22.1 | Standard ICLcont | 53.1 | 55.6 | 67.9 |
Zero-shotstop | 40.6 | 38.8 | 49.1 | Standard ICLstop | 57.5 | 59.8 | 72.0 |
Zero-shottemp | 51.1 | 53.7 | 67.6 | Standard ICLtemp | 43.3 | 42.6 | 53.9 |
Llama 7B | |||||||
Setting | Ins. 1 | Ins. 2 | Ins. 3 | Setting | Ins. 1 | Ins. 2 | Ins. 3 |
Zero-shotcont | 31.1 | 30.2 | 35.2 | Standard ICLcont | 70.6 | 78.2 | 75.2 |
Zero-shotstop | 51.4 | 63.5 | 61.8 | Standard ICLstop | 73.9 | 80.1 | 79.4 |
Zero-shottemp | 62.5 | 73.2 | 71.1 | Standard ICLtemp | 59.9 | 69.1 | 74.0 |
OpenLlama 3B | |||||||
Setting | Ins. 1 | Ins. 2 | Ins. 3 | Setting | Ins. 1 | Ins. 2 | Ins. 3 |
Zero-shotcont | 6.7 | 6.3 | 7.2 | Standard ICLcont | 56.1 | 60.4 | 58.1 |
Zero-shotstop | 43.1 | 48.3 | 40.2 | Standard ICLstop | 58.1 | 61.4 | 59.6 |
Zero-shottemp | 55.7 | 59.9 | 57.7 | Standard ICLtemp | 48.6 | 54.0 | 47.8 |
Llama 7B | |||||||
Setting | Ins. 1 | Ins. 2 | Ins. 3 | Setting | Ins. 1 | Ins. 2 | Ins. 3 |
Zero-shotcont | 15.0 | 15.9 | 6.8 | Standard ICLcont | 64.9 | 66.1 | 68.7 |
Zero-shotstop | 49.7 | 48.1 | 48.6 | Standard ICLstop | 66.8 | 67.6 | 69.6 |
Zero-shottemp | 66.1 | 66.5 | 69.0 | Standard ICLtemp | 58.9 | 59.2 | 61.7 |
OpenLlama 3B | |||||||
Settings | En-Fr | En-De | En-Da | Settings | En-Fr | En-De | En-Da |
Zero-shotcont | 0.13 | 0.38 | 0.28 | Standard ICLcont | 26.07 | 12.53 | 17.43 |
Zero-shotstop | 16.68 | 9.17 | 13.09 | Standard ICLstop | 26.19 | 12.52 | 17.29 |
Zero-shottemp | 26.06 | 12.92 | 17.17 | Standard ICLtemp | 17.38 | 8.88 | 12.9 |
Llama 7B | |||||||
Settings | En-Fr | En-De | En-Da | Settings | En-Fr | En-De | En-Da |
Zero-shotcont | 11.76 | 13.83 | 10.18 | Standard ICLcont | 35.39 | 24.23 | 30.08 |
Zero-shotstop | 30.23 | 21.76 | 23.34 | Standard ICLstop | 35.36 | 24.33 | 29.99 |
Zero-shottemp | 35.47 | 24.34 | 30.12 | Standard ICLtemp | 31.09 | 21.98 | 24.88 |
Appendix L Representation-level ablation on Machine Translation Tasks
Besides the classification tasks, we also show results in the machine translation tasks to show that our findings could also be extended in text generation tasks. We used the Flores MT dataset (Costa-jussà et al., 2022) to conduct this set of 4-shot machine translation experiments. The results are reported with the BLEU metric (Papineni et al., 2002). We investigated three different language directions: English-to-French, English-to-Danish, and English-to-German. We used 10 random seeds for En-Fr and En-De and 15 random seeds for En-Da to randomly choose the demonstrations. 100 test examples are sampled in this set of the experiments as a computational compromise. Similar to the classification tasks, we keep the answer (i.e., target language) unablated for all the settings and ablate different kinds of tokens. Results in Table 23 show the consistent finding to those in Section 5.2.1.
Appendix M Representation-level ablation on Question Answering Tasks
We show the representation-level experimental results of the question answering (QA) tasks in this section. We used Commonsense QA (Talmor et al., 2019) dataset to test if the template and stopword tokens would directly affect the downstream task performance. We applied the settings of 4 in-context examples and 15 random seeds in this set of experiments. We frame the task as directly answering the questions instead of choosing one answer from the choices because the token types in this scenario are easier to be categorized.
Results shown in Table 24 demonstrate that our main findings, that template and stopword tokens are more likely to serve as performance-critical tokens, still hold in the QA tasks.
Setting | OpenLlama 3B | Llama 7B | Llama 13B | Llama 33B | Llama 2 7B | Llama 2 13B | Mistral 7B |
Zero-shotcont | 7.42 | 16.62 | 14.49 | 19.47 | 17.51 | 17.78 | 16.64 |
Zero-shotstop | 11.71 | 21.96 | 18.98 | 23.38 | 22.13 | 22.31 | 19.16 |
Zero-shottemp | 13.24 | 24.38 | 25.73 | 27.20 | 25.42 | 25.11 | 25.56 |
Standard ICLcont | 14.40 | 24.07 | 26.22 | 27.73 | 25.89 | 24.71 | 25.47 |
Standard ICLstop | 11.96 | 21.84 | 21.44 | 26.93 | 24.62 | 23.09 | 23.69 |
Standard ICLtemp | 6.89 | 16.29 | 15.31 | 19.78 | 18.51 | 16.64 | 15.51 |
Appendix N Representation-level Ablation based on the Token Count
One possible explanation for the performance variation when different types of tokens are ablated at the representation level is the simple fact that the number of tokens being ablated may vary. Intuitively, template and stopword tokens are far fewer in number compared to content tokens. In this section, we show the statistics of the number count of each type of tokens and include a supplementary experiment that only let the LLM attend to certain number of token representations of each type of the tokens.
We first present the average token count for each type of token across the datasets. Token counts may vary depending on the tokenizer used by the large language models, and all statistics are shown in Table 25. The results indicate that the number of template and stopword tokens is much smaller than the number of content tokens, suggesting that performance variation during ablation is not solely due to differences in token type counts.
Tokenizer | Setting | Avg. | AGNews | DBPedia | SST2 | TREC | CB | RTE |
OpenLlama 3B | cont | 204.1 | 207.5 | 278.8 | 116.3 | 48.7 | 295.5 | 278.0 |
stop | 43.0 | 43.2 | 45.1 | 27.7 | 21.8 | 66.7 | 53.7 | |
temp | 56.5 | 43.3 | 49.9 | 42.4 | 48.8 | 78.5 | 76.3 | |
Llama & Llama 2 | cont | 220.6 | 238.5 | 288.8 | 127.8 | 51.3 | 312.0 | 305.1 |
stop | 45.0 | 43.1 | 46.1 | 29.9 | 21.7 | 71.1 | 57.9 | |
temp | 52.7 | 41.3 | 50.5 | 48.7 | 36.7 | 70.9 | 68.1 | |
Mistral | cont | 213.6 | 225.0 | 284.7 | 123.7 | 50.1 | 304.2 | 293.7 |
stop | 44.7 | 43.2 | 45.0 | 29.9 | 21.7 | 70.7 | 57.8 | |
temp | 54.6 | 45.3 | 53.5 | 40.5 | 40.7 | 75.0 | 72.5 |
We then conduct an additional ablation experiment in which the model attends to representations from a specific number of tokens of a given type. We also include a baseline where a random subset of token representations from all prompt tokens is unmasked to the test examples. In this set of experiments, the label tokens are always included and are not counted as part of the token numbers.
Results in Table 26 demonstrate that when the model is exposed to an equal number of each type of token representation, the performance consistently improves with template and stopword tokens, outperforming both content tokens and the random baseline. In contrast, models attending to the same number of content tokens consistently underperform relative to the random baseline. Additionally, all results improve when more tokens are included in the attention of test examples. This experiment further supports our claim that template and stopword tokens are more likely to serve as performance-critical tokens.
Models | Setting | AGNews | DBPedia | SST2 | TREC | CB | RTE | Avg. |
10 Random Tokens | ||||||||
Llama 2 7B | From all | 14.5 | 16.9 | 71.7 | 47.3 | 60.2 | 57.5 | 44.7 |
From cont | 3.0 | 9.2 | 60.9 | 61.9 | 65.2 | 60.4 | 43.4 | |
From stop | 15.2 | 35.6 | 75.2 | 68.5 | 62.0 | 64.1 | 53.4 | |
From temp | 53.0 | 50.1 | 59.1 | 56.2 | 64.4 | 60.2 | 57.2 | |
Mistral 7B | From all | 67.8 | 67.2 | 75.0 | 70.9 | 63.8 | 60.7 | 67.6 |
From cont | 36.6 | 64.0 | 63.1 | 67.5 | 59.3 | 57.3 | 58.0 | |
From stop | 73.5 | 68.6 | 86.3 | 72.8 | 61.8 | 61.6 | 70.8 | |
From temp | 78.7 | 78.7 | 92.1 | 68.9 | 71.2 | 68.4 | 76.3 | |
20 Random Tokens | ||||||||
Llama 2 7B | From all | 12.6 | 24.0 | 77.2 | 54.4 | 60.6 | 58.9 | 47.9 |
From cont | 1.3 | 8.6 | 65.3 | 63.5 | 61.7 | 62.2 | 43.8 | |
From stop | 28.3 | 45.5 | 84.5 | 66.8 | 58.0 | 67.1 | 58.4 | |
From temp | 75.0 | 63.7 | 60.3 | 59.0 | 65.4 | 60.7 | 64.0 | |
Mistral 7B | From all | 78.4 | 67.8 | 74.0 | 70.1 | 65.4 | 62.0 | 69.6 |
From cont | 69.1 | 63.3 | 59.3 | 68.2 | 60.5 | 56.8 | 62.9 | |
From stop | 78.4 | 74.3 | 89.9 | 74.2 | 68.3 | 69.1 | 75.7 | |
From temp | 81.9 | 78.7 | 93.0 | 68.9 | 72.4 | 71.4 | 77.7 | |
30 Random Tokens | ||||||||
Llama 2 7B | From all | 27.4 | 27.7 | 77.9 | 54.9 | 63.9 | 62.2 | 52.3 |
From cont | 1.3 | 7.6 | 77.9 | 66.3 | 64.2 | 63.5 | 46.8 | |
From stop | 44.8 | 59.8 | 92.8 | 68.3 | 53.2 | 69.3 | 64.7 | |
From temp | 76.9 | 66.1 | 61.0 | 59.6 | 67.9 | 59.8 | 65.2 | |
Mistral 7B | From all | 79.9 | 74.6 | 83.0 | 67.8 | 69.0 | 64.6 | 73.2 |
From cont | 71.0 | 62.6 | 58.2 | 67.2 | 59.2 | 59.4 | 62.9 | |
From stop | 79.8 | 76.9 | 91.0 | 74.2 | 71.1 | 68.8 | 77.0 | |
From temp | 84.0 | 79.3 | 93.8 | 68.9 | 75.6 | 75.4 | 79.5 | |
40 Random Tokens | ||||||||
Llama 2 7B | From all | 45.2 | 27.2 | 77.9 | 54.9 | 58.9 | 62.3 | 54.4 |
From cont | 0.9 | 9.1 | 89.5 | 66.1 | 62.3 | 62.4 | 48.4 | |
From stop | 49.2 | 63.7 | 94.0 | 68.7 | 53.7 | 70.8 | 66.7 | |
From temp | 77.9 | 66.1 | 62.9 | 61.5 | 68.1 | 60.1 | 66.1 | |
Mistral 7B | From all | 78.8 | 73.4 | 86.9 | 69.2 | 69.3 | 67.4 | 74.2 |
From cont | 74.1 | 59.4 | 55.8 | 68.0 | 61.0 | 59.9 | 63.0 | |
From stop | 80.1 | 76.4 | 91.0 | 74.2 | 72.0 | 69.6 | 77.2 | |
From temp | 85.0 | 80.7 | 93.8 | 69.0 | 76.4 | 76.2 | 80.2 |
Appendix O Results of the Token-level Ablation
Detailed results of the token-level ablation are shown in Table 27. We omited the temp case from here since it constantly yields an accuracy of 0% when both the template token in the demonstrations and the test examples are ablated. Since the setting for the template tokens are not aligned with the ones for the stopword and content tokens, we included another set of experiments where only the the template tokens in the demonstrations are ablated at the token level in Appendix P. We want to emphasize that this experimental design choice does not affect the main findings in Section 5.2.2, where information is being propagated from the representations of content tokens to the representations of the performance-critical tokens and this incorporation of the information makes them better at affecting task performance, partially explaining the working mechanism of in-context learning.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
cont | 31.5 | 63.0 | 40.6 | 25.4 | 56.1 | 48.9 | 44.3 | |
stop | 64.4 | 91.5 | 20.9 | 62.3 | 57.8 | 52.6 | 58.3 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
cont | 55.2 | 67.2 | 42.6 | 50.8 | 57.4 | 56.3 | 54.9 | |
stop | 82.3 | 93.8 | 64.1 | 69.7 | 66.5 | 70.0 | 74.4 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
cont | 78.8 | 81.7 | 45.3 | 75.1 | 55.1 | 54.5 | 65.1 | |
stop | 82.5 | 92.5 | 61.5 | 76.5 | 69.6 | 40.5 | 70.5 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
cont | 74.0 | 89.6 | 67.0 | 73.0 | 69.8 | 49.0 | 70.4 | |
stop | 85.3 | 96.4 | 66.9 | 77.9 | 77.7 | 81.3 | 80.9 |
Appendix P Token-level Ablation for Template Tokens
To maintain consistency of the templates across both demonstrations and test examples, we choose to ablate the template tokens at the token level in both in Section 5.2.2. This experimental design differs from the other two token-level ablations. This inconsistency does not impact the main findings in Section 5.2.2, which show that information is propagated from the representations of content tokens to the representations of performance-critical tokens and this information aggregation enhances the ability of performance-critical tokens to improve the final task performance, partially explaining the mechanism of in-context learning. For completeness, we provide a supplemental experiment in this section where only the template tokens in the demonstrations are ablated.
Results in Table 28 demonstrate that, although not all values reduce to 0%, large language models perform significantly worse than in the standard in-context learning case and the other two ablation scenarios after the removal of template tokens from the demonstrations except for a few rare cases. This further supports the finding that template tokens are likely important as performance-critical tokens.
Models | Settings | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
OpenLlama 3B | Standard ICL | 58.0 | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 |
temp | 17.4 | 0.0 | 24.2 | 40.5 | 0.4 | 35.6 | 3.6 | |
Llama 7B | Standard ICL | 74.8 | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 |
temp | 29.0 | 41.0 | 11.4 | 39.9 | 0.9 | 49.7 | 31.3 | |
Llama 13B | Standard ICL | 70.4 | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 |
temp | 36.9 | 82.1 | 17.5 | 51.4 | 8.6 | 39.5 | 22.0 | |
Llama 33B | Standard ICL | 81.6 | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 |
temp | 63.0 | 73.5 | 82.8 | 58.8 | 39.9 | 66.4 | 56.7 |
Appendix Q Possible Applications
In this section, we discuss several potential applications that could benefit from the findings in our work. These include long sequence processing, where our insights can help models handle longer contexts more efficiently; in-context learning with more demonstrations, enabling the inclusion of additional examples without compromising performance; better ICL prompt designing and engineering, improving the creation of more effective prompts; and improving model robustness, ensuring consistent performance despite prompt variations. Each of these areas can be enhanced by understanding the role of performance-critical tokens in large language models.
Long sequence processing
As discussed in our paper, we hypothesize that stopword tokens tend to function as performance-critical tokens by encapsulating the semantics of preceding tokens. This finding suggests an opportunity to improve the efficiency of modeling longer sequences by selectively deleting or compressing certain hidden states during the encoding and generation stages of large language models (LLMs). Specifically, by retaining only the essential performance-critical representations while reducing unnecessary content from less informative tokens, models could manage longer inputs and outputs without compromising performance. This approach not only conserves computational resources but also addresses token length limitations in LLMs, allowing for extended sequence processing and potentially more nuanced learning from longer contexts.
This area is indeed attracting increased research attention, and our findings could contribute valuable insights into ongoing work on efficient sequence modeling and memory management in LLMs (Liu et al., 2023b; Zhang et al., 2023; Bai et al., 2024). By identifying which tokens retain critical task-related information, our work aligns with and can inform methods focused on compressing intermediate states and improving long-context processing for applications ranging from summarization and document understanding to interactive dialogue systems.
In-context learning with more demonstrations
Given our findings, there is a promising avenue for improving in-context learning (ICL) performance by including a greater number of examples in ICL prompts (Li et al., 2023a; Hao et al., 2022; Bertsch et al., 2024). Our results suggest that only a subset of token representations, specifically performance-critical tokens, play a critical role in determining ICL performance, while the representations of other tokens are less impactful. This observation opens up the possibility of selectively compressing or omitting unimportant token representations after the initial encoding of a demonstration. By doing so, it becomes feasible to maximize the use of the model’s fixed-length capacity, potentially enabling the inclusion of a higher number of examples within the same prompt length constraints. This approach may enhance the effectiveness of ICL in tasks where the availability of diverse examples contributes to improved model accuracy and stability.
Better ICL prompt designing and engineering
Our investigation into which components of ICL prompts are most critical for task performance is worthwhile and useful for directing where to put effort into tuning or improving prompts. Furthermore, the exploration on the characteristics of performance-critical tokens are useful for future design choices in ICL prompting, and help the field understand why some prompts work better than others for ICL. For instance, knowing that template and stopword tokens are particularly performance-critical allows developers to optimize prompts by focusing on specific token structures or repetitions that are most influential. This insight can improve task performance consistency across variations in prompt phrasing and structure, ultimately making prompt creation more efficient and predictable.
Improving Model Robustness
The findings in our study can also inform techniques to enhance the robustness of large language models (LLMs). Since prompt sensitivity (e.g., to token arrangement) can often lead to fluctuations in performance, understanding performance-critical tokens helps mitigate these vulnerabilities. By aligning model training and prompt engineering to leverage performance-critical token characteristics, it becomes possible to minimize performance drops due to minor prompt alterations, thereby enhancing the stability and reliability of LLMs in production environments.
Appendix R Results of Representation-level Partial performance-critical Token Ablation
The full results on all the six datasets are shown in Table 29 and Table 30. Most of the results align with our descriptions in Section C, where the performance-critical tokens should be utilized together to provide the best performance and that removing some of them would cause performance degeneration, demonstrated by the performance decrease from Table 29, or instability issues, shown by Table 30.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||
OpenLlama 3B | temp with | 56.5 | 47.4 | 86.7 | 83.7 | 27.1 | 26.5 | 62.2 | 59.8 | 56.4 | 56.0 | 52.3 | 56.1 |
50.3 | 47.1 | 85.7 | 84.4 | 28.9 | 24.4 | 57.7 | 57.7 | 56.5 | 56.1 | 53.2 | 55.2 | ||
34.6 | 32.7 | 86.9 | 82.3 | 28.2 | 31.2 | 55.5 | 54.1 | 58.3 | 59.2 | 55.4 | 58.3 | ||
Llama 7B | temp with | 70.8 | 57.3 | 90.2 | 87.1 | 58.4 | 46.7 | 66.2 | 63.8 | 66.3 | 59.5 | 73.5 | 69.6 |
62.7 | 55.1 | 91.6 | 87.1 | 52.8 | 43.3 | 61.6 | 61.8 | 58.9 | 56.5 | 59.2 | 55.7 | ||
50.8 | 48.6 | 84.9 | 82.8 | 46.0 | 50.2 | 57.9 | 55.2 | 56.7 | 56.7 | 66.2 | 64.5 | ||
Llama 13B | temp with | 80.0 | 76.2 | 92.3 | 89.1 | 58.6 | 54.0 | 76.9 | 71.4 | 68.5 | 59.8 | 47.7 | 35.0 |
79.9 | 76.3 | 91.5 | 88.9 | 55.1 | 47.8 | 75.8 | 70.7 | 65.5 | 59.0 | 35.7 | 24.5 | ||
72.0 | 72.1 | 81.1 | 75.9 | 47.1 | 48.3 | 60.3 | 35.5 | 61.2 | 58.4 | 36.2 | 36.0 | ||
Llama 33B | temp with | 80.5 | 75.0 | 95.2 | 93.3 | 65.2 | 66.7 | 75.2 | 73.5 | 79.0 | 77.1 | 80.0 | 70.7 |
78.7 | 71.5 | 95.2 | 92.8 | 68.1 | 67.7 | 75.1 | 73.8 | 77.4 | 75.4 | 73.3 | 62.3 | ||
69.2 | 69.5 | 93.9 | 92.9 | 62.1 | 66.2 | 71.3 | 70.1 | 72.8 | 70.0 | 67.4 | 63.5 |
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | with “:” | w/o “:” | ||
OpenLlama 3B | temp w/o | 41.5 | 54.6 | 14.3 | 73.2 | 36.5 | 42.0 | 29.4 | 21.7 | 24.7 | 45.7 | 0.7 | 3.5 |
42.2 | 52.2 | 18.5 | 79.9 | 39.7 | 42.2 | 22.6 | 22.5 | 49.8 | 57.1 | 3.1 | 6.5 | ||
36.3 | 35.5 | 83.0 | 83.4 | 43.2 | 41.9 | 16.3 | 18.7 | 54.4 | 56.8 | 1.2 | 4.2 | ||
Llama 7B | temp w/o | 50.4 | 56.6 | 68.2 | 56.1 | 55.3 | 48.5 | 0.2 | 1.3 | 40.7 | 42.5 | 28.5 | 18.8 |
46.1 | 50.6 | 61.9 | 55.1 | 43.5 | 44.7 | 0.0 | 0.2 | 43.7 | 49.9 | 27.1 | 26.5 | ||
21.4 | 12.7 | 86.2 | 66.5 | 54.4 | 55.6 | 0.0 | 0.0 | 56.0 | 55.6 | 39.4 | 35.1 | ||
Llama 13B | temp w/o | 66.9 | 77.0 | 65.6 | 87.9 | 51.8 | 53.1 | 0.1 | 0.1 | 57.5 | 53.7 | 16.7 | 21.9 |
72.9 | 76.6 | 83.0 | 89.5 | 45.5 | 48.4 | 0.0 | 0.0 | 53.6 | 52.8 | 16.0 | 20.1 | ||
79.2 | 77.7 | 77.5 | 47.5 | 56.8 | 43.2 | 0.0 | 0.0 | 54.8 | 53.7 | 4.3 | 2.4 | ||
Llama 33B | temp w/o | 77.3 | 78.2 | 17.3 | 88.9 | 65.4 | 69.3 | 31.0 | 41.7 | 71.8 | 65.8 | 23.8 | 23.0 |
72.9 | 72.4 | 29.2 | 87.4 | 65.6 | 70.9 | 14.9 | 37.9 | 70.6 | 67.8 | 19.9 | 21.1 | ||
69.5 | 74.3 | 92.6 | 92.8 | 70.0 | 70.8 | 42.0 | 20.3 | 67.0 | 61.3 | 23.1 | 18.5 |
Appendix S Significance test for the representation-level ablation
In this section, we report the p-value of all the pair-wise comparisons in the representation-level ablation experiments in Table 2 and Table 3. Results are shown in Table 31. Most of the ablation results show significant difference among different ablation scenarios.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||||||
P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | P-value | p < 0.05 | ||
OpenLlama 3B | temp <-> cont | 0.0000581 | T | 0.0000000 | T | 0.1952165 | F | 0.0000000 | T | 0.0042427 | T | 0.0027293 | T |
temp <-> stop | 0.0001605 | T | 0.0571278 | F | 0.0242797 | T | 0.0000663 | T | 0.0319815 | T | 0.0985942 | F | |
cont <-> stop | 0.0023957 | T | 0.0000001 | T | 0.1940792 | F | 0.0000000 | T | 0.0000073 | T | 0.0000698 | T | |
temp_cont <-> cont_stop | 0.0000065 | T | 0.0385760 | T | 0.3206221 | F | 0.0000000 | T | 0.1237570 | F | 0.0544049 | F | |
temp_stop <-> cont_stop | 0.0001166 | T | 0.4514005 | F | 0.2549225 | F | 0.0000001 | T | 0.0545474 | F | 0.0534521 | F | |
temp_cont <-> temp_stop | 0.0000507 | T | 0.0096752 | T | 0.0005775 | T | 0.0000000 | T | 0.1208140 | F | 0.3193696 | F | |
Llama 7B | temp <-> cont | 0.0000000 | T | 0.0000001 | T | 0.0000020 | T | 0.0000001 | T | 0.0000004 | T | 0.0000001 | T |
temp <-> stop | 0.0000083 | T | 0.0101283 | T | 0.0002883 | T | 0.1438193 | F | 0.0000000 | T | 0.0000031 | T | |
cont <-> stop | 0.0000060 | T | 0.0000001 | T | 0.0019529 | T | 0.0000001 | T | 0.3392237 | F | 0.0016487 | T | |
temp_cont <-> cont_stop | 0.0000115 | T | 0.0030175 | T | 0.0005950 | T | 0.0000000 | T | 0.0000000 | T | 0.0001649 | T | |
temp_stop <-> cont_stop | 0.0002004 | T | 0.0227328 | T | 0.0015468 | T | 0.0000000 | T | 0.0000001 | T | 0.0000094 | T | |
temp_cont <-> temp_stop | 0.0086396 | T | 0.0003632 | T | 0.0089932 | T | 0.0000007 | T | 0.1637608 | F | 0.1553081 | F | |
Llama 13B | temp <-> cont | 0.0000000 | T | 0.0000000 | T | 0.0000082 | T | 0.0000001 | T | 0.0006445 | T | 0.1060226 | F |
temp <-> stop | 0.0003841 | T | 0.0000012 | T | 0.0034370 | T | 0.0002018 | T | 0.0000000 | T | 0.0010178 | T | |
cont <-> stop | 0.0000000 | T | 0.0000202 | T | 0.0002820 | T | 0.0000000 | T | 0.0098209 | T | 0.0022848 | T | |
temp_cont <-> cont_stop | 0.0010838 | T | 0.0000730 | T | 0.0004557 | T | 0.0000048 | T | 0.0000001 | T | 0.0002364 | T | |
temp_stop <-> cont_stop | 0.0007763 | T | 0.0000310 | T | 0.0016544 | T | 0.0000000 | T | 0.0000000 | T | 0.0000888 | T | |
temp_cont <-> temp_stop | 0.4411518 | F | 0.1158895 | F | 0.3323328 | F | 0.0000000 | T | 0.3148253 | F | 0.0144961 | T | |
Llama 33B | temp <-> cont | 0.0000000 | T | 0.0000003 | T | 0.1534319 | F | 0.0000000 | T | 0.0000000 | T | 0.0002244 | T |
temp <-> stop | 0.0007359 | T | 0.0048547 | T | 0.1797405 | F | 0.0000023 | T | 0.0000002 | T | 0.0008789 | T | |
cont <-> stop | 0.0000000 | T | 0.0000003 | T | 0.0204911 | T | 0.0000000 | T | 0.0002626 | T | 0.4319440 | F | |
temp_cont <-> cont_stop | 0.0001365 | T | 0.0788756 | F | 0.0032131 | T | 0.0000000 | T | 0.0000098 | T | 0.0003242 | T | |
temp_stop <-> cont_stop | 0.0006045 | T | 0.0609501 | F | 0.0165374 | T | 0.0000011 | T | 0.0000009 | T | 0.0003821 | T | |
temp_cont <-> temp_stop | 0.0012936 | T | 0.3583931 | F | 0.1415489 | F | 0.0001034 | T | 0.0009055 | T | 0.3979685 | F |
Appendix T Template used for the random string experiments
In this section, we present all the in-context learning templates used for the random experiments in Section 6. In the Randomfixed scenario, the and are consistent across all demonstrations. For the Randomnonfixed scenario, we employ different random string templates for each demonstration. We use 5 random string templates for each setting, shown in Table 37, Table 38, Table 39, Table 40, and Table 41. The results in Section 6 are averaged over the results with all the different random string templates.
Appendix U Supplemental Experiments for the Repetition Characteristic
In Section 6, we examine the repetition characteristic of performance-critical tokens with random template tokens, which could not be general enough since random string tokens are less used in real-world applications. Hence, we conduct another set of experiments in this section, using template tokens with lexical meanings to test the characteristic of repetition.
These experiments includes two sets of comparisons shown in Table 32 and Table 33. The first set of templates uses meaningful, normal words but exhibits less lexical similarity to the task. The second set of templates is more closely related to the task. All comparisons are made between non-repetitive and repetitive cases.
The results presented in Table 34 show that, when random strings without lexical meanings are not used, the repetitive patterns can also enhance the final performances and help encode the task within the representations of template tokens, proving our claim that repetition is an important characteristic of performance-critical tokens.
Template 1 |
Classify the news articles into the categories of World, Sports, Business, and Technology. |
dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
cat: Technology |
juice: Amazon’s Holiday Pi. Leave it to Amazon.com (Nasdaq: AMZN). Apparently, the holiday season could be a rich opportunity to addict more users to Amazon’s A9. |
wine: Technology |
sleep: Will historic flight launch space tourism?. Regardless, space competitions are poised to become big business. |
wake: Technology |
bunny: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe … |
easter: |
Template 2 |
Classify the news articles into the categories of World, Sports, Business, and Technology. |
dog: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
cat: Technology |
dog: Amazon’s Holiday Pi. Leave it to Amazon.com (Nasdaq: AMZN). Apparently, the holiday season could be a rich opportunity to addict more users to Amazon’s A9. |
cat: Technology |
dog: Will historic flight launch space tourism?. Regardless, space competitions are poised to become big business. |
cat: Technology |
dog: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
cat: |
Template 3 |
Classify the news articles into the categories of World, Sports, Business, and Technology. |
article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
answer: Technology |
input: Amazon’s Holiday Pi. Leave it to Amazon.com (Nasdaq: AMZN). Apparently, the holiday season could be a rich opportunity to addict more users to Amazon’s A9. |
output: Technology |
text: Will historic flight launch space tourism?. Regardless, space competitions are poised to become big business. |
label: Technology |
sentence: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
result: |
Template 4 |
Classify the news articles into the categories of World, Sports, Business, and Technology. |
article: First class to the moon. London - British airline magnate Richard Branson announced a plan on Monday for the world’s first commercial space flights, saying "thousands" of fee-paying astronauts could be sent into orbit in the near future. |
answer: Technology |
article: Amazon’s Holiday Pi. Leave it to Amazon.com (Nasdaq: AMZN). Apparently, the holiday season could be a rich opportunity to addict more users to Amazon’s A9. |
answer: Technology |
article: Will historic flight launch space tourism?. Regardless, space competitions are poised to become big business. |
answer: Technology |
article: SMART-1 makes lunar orbit. The SMART-1 probe has entered its lunar orbit, and the history books as the first European mission to have done so. Professor David Southwood, director of science for the European Space Agency (ESA), said: "Europe… |
answer: |
Models | Setting | AGNews | DBPedia | TREC | Avg. |
OpenLlama 3B | Template 1 | 33.56 | 9.72 | 5.84 | 16.37 |
Template 2 | 55.56 | 61.44 | 22.36 | 46.45 | |
Template 3 | 48.24 | 50.16 | 24.84 | 41.08 | |
Template 4 | 69.64 | 63.20 | 20.96 | 51.27 | |
Llama 7B | Template 1 | 6.00 | 0.24 | 12.92 | 6.39 |
Template 2 | 26.52 | 51.20 | 25.32 | 34.35 | |
Template 3 | 19.80 | 62.28 | 1.88 | 27.99 | |
Template 4 | 36.44 | 64.68 | 18.68 | 39.93 | |
Llama 13B | Template 1 | 7.40 | 0.00 | 5.68 | 4.36 |
Template 2 | 46.68 | 71.40 | 35.80 | 51.29 | |
Template 3 | 15.60 | 76.36 | 4.96 | 32.31 | |
Template 4 | 49.08 | 75.56 | 23.80 | 49.48 | |
Llama 2 7B | Template 1 | 52.84 | 12.08 | 35.60 | 33.51 |
Template 2 | 75.60 | 82.80 | 56.04 | 71.48 | |
Template 3 | 32.96 | 76.56 | 7.04 | 38.85 | |
Template 4 | 70.16 | 80.96 | 58.24 | 69.79 | |
Llama 2 13B | Template 1 | 8.28 | 0.04 | 2.68 | 3.67 |
Template 2 | 19.80 | 44.52 | 13.92 | 26.08 | |
Template 3 | 5.84 | 59.88 | 1.72 | 22.48 | |
Template 4 | 28.60 | 65.04 | 12.64 | 35.43 | |
Mistral 7B | Template 1 | 27.68 | 0.20 | 17.96 | 15.28 |
Template 2 | 67.48 | 67.60 | 31.20 | 55.43 | |
Template 3 | 2.64 | 47.68 | 4.04 | 18.12 | |
Template 4 | 59.12 | 70.64 | 39.12 | 56.29 |
Appendix V Supplemental Experiments for the Structural Cue Characteristic
In this section, we describe a set of supplemental experiments, which support the characteristic of structural cues from the perspective of representation-level ablation. An intuitive method to verify the effect of the structural cues would be using the same random strings to replace and , making it harder for a model to parse the structure of the text. However, this would bring the factor of repetition into the process, potentially confounding the results. Hence, we instead design a one-shot Randomfixed experiment. The one-shot Randomfixed setting allows us to control both the characteristics of lexical meaning and repetition since the templates are made up of random strings and there is only one training demonstration. With these two characteristics controlled, we use the masking ablation method from Section 5.2.1 to confirm to what extent these random string tokens can function effectively as delimiters between inputs and outputs in ICL prompts. Specifically, we include results from the Zero-shot temp and Zero-shot “:” scenarios, as well as the standard results of one-shot Randomfixed, for a more comprehensive analysis, shown in Table 35. Examples of all the different model variants are shown in Appendix W.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | One-shot Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 |
Zero-shottemp | 39.5 | 49.8 | 27.7 | 13.3 | 49.8 | 44.9 | 37.5 | |
Zero-shot“:” | 31.5 | 35.9 | 23.8 | 8.0 | 35.9 | 33.8 | 28.2 | |
Llama 7B | One-shot Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 |
Zero-shottemp | 2.1 | 15.5 | 7.6 | 3.7 | 15.5 | 5.4 | 8.3 | |
Zero-shot“:” | 3.6 | 7.5 | 14.6 | 3.0 | 7.5 | 6.8 | 7.2 | |
Llama 13B | One-shot Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 |
Zero-shottemp | 29.2 | 48.9 | 36.1 | 35.7 | 48.9 | 14.0 | 35.5 | |
Zero-shot“:” | 14.3 | 22.4 | 25.4 | 22.5 | 22.4 | 28.9 | 22.7 | |
Llama 33B | One-shot Randomfixed | 69.7 | 53.0 | 37.8 | 72.8 | 53.0 | 37.6 | 54.0 |
Zero-shottemp | 61.2 | 56.3 | 41.1 | 69.2 | 56.3 | 43.0 | 54.5 | |
Zero-shot“:” | 43.3 | 41.8 | 37.4 | 65.0 | 41.8 | 39.5 | 44.8 |
We observe that adding the attention to random template token representations in the one-shot setting often leads to performance increases while masking the attention to the template tokens and only attending to “:” leads to performance decreases. This indicates that the presence of these tokens is critical to maintaining task performance. With all other characteristics being controlled, this leads us to believe that the delimiting nature of template tokens is likely to be an important characteristic in their role as performance-critical tokens.
Appendix W Discussion about the Characteristic of Structural Cue
As discussed in Section 6, we view structural cues as the textual and structural cues present in the prompt allowing the model to distinguish between the different parts of the ICL demonstration. We believe that performance-critical tokens naturally play this role since the same types of tokens are likely to delimit pretraining text (e.g., html, markdown, etc.). An example of how we believe performance-critical tokens naturally delimit an ICL prompt is shown in Table 36, sampled from the SST2 dataset tested in our experiments. We bold and place in brackets the role of each section of the prompt as well as what types of tokens it contains.
During LLM pre-training, it is very likely that the model has seen text formatted in a similar way (e.g. in html, plain text with headings), in which the LLM would learn to recognize and store information in the representation of these formatting tokens. A recent study also provides supporting facts from the perspective of pretraining for this (Chen et al., 2024).
[…]
Zero-shot temp and Zero-shot “:” To investigate whether these random string tokens are working as performance-critical tokens, given that they only serve as providing structural cues, we applied the representation-level ablation to see the model’s performance when the test examples have or do not have access to the representations of these random string tokens, comparing the performance among [one-shot Randomfixed], [Zero-shot temp] and [Zero-shot “:”]. [Zero-shot temp] and [Zero-shot “:”] are all the representation-level ablation models based on one-shot Randomfixed, where the templates in these settings are all random strings, shown in Table 36.
The results in Table 35 show that in this setting, the model could still store the task-related information in the representations of the random string tokens, shown by the performance drop when removing their representations. There is nothing else for the model to recognize these random strings and store the information in their representations except that these tokens serve as delimiters to inform the model distinguishing the different parts of the prompt.
Standard ICL |
Classify the reviews into the categories of Positive and Negative. [instruction] |
Review: [delimiter: template] |
Peppered with witty dialogue and inventive moments. [demonstration: content + stopword] |
Answer: [delimiter: template] |
Positive [label] |
One-shot Randomfixed |
Classify the reviews into the categories of Positive and Negative. [instruction] |
dsafjkldafdsajk: [delimiter: random template 1] |
Peppered with witty dialogue and inventive moments. [demonstration] |
reqwiorewsdafjl: [delimiter: random template 2] |
Positive [label] |
Zero-shottemp |
Classify the reviews into the categories of Positive and Negative. [instruction] |
dsafjkldafdsajk: [delimiter: random template 1] |
<m><m><m> … <m> [masked demonstration] |
reqwiorewsdafjl: [delimiter: random template 2] |
Positive [label] |
Zero-shot“:” |
Classify the reviews into the categories of Positive and Negative. [instruction] |
<m>: [delimiter: random template 1] |
<m><m><m> … <m> [masked demonstration] |
<m>: [delimiter: random template 2] |
Positive [label] |
Datasets | Notations | Examples |
Randomfixed | ||
[…]
Randomnonfixed | ||
[…]
Swap | ||
[…]
Datasets | Notations | Examples |
Randomfixed | ||
[…]
Randomnonfixed | ||
[…]
Datasets | Notations | Examples |
Randomfixed | ||
[…]
Randomnonfixed | ||
[…]
Datasets | Notations | Examples |
Randomfixed | ||
[…]
Randomnonfixed | ||
[…]
Datasets | Notations | Examples |
Randomfixed | ||
[…]
Randomnonfixed | ||
[…]